OpenStudy (anonymous):

Topic: Base e and Natural Logarithms ln (x-2) = 2 and another one is ln x + ln 2x = 2

OpenStudy (johnweldon1993):

Hint* \[\huge e^{ln(x - 2) = (x - 2)}\]

OpenStudy (anonymous):

\[x-2=e^2,x=?\]

OpenStudy (anonymous):

I'm confused :(

OpenStudy (anonymous):

\[\ln a+\ln b=\ln ab\]

OpenStudy (anonymous):

so ln 2x^2

OpenStudy (johnweldon1993):

Well lets tackle question 1...you said you were confused....what on?

OpenStudy (anonymous):

all of it because I thought e was ln

OpenStudy (johnweldon1993):

e is the inverse of ln Hence...when you raise e^ln power....they cancel out...that is key in logarithm problems \[\large ln(x - 2) = 2\] raise e to both sides \[\huge e^{ln(x - 2)} = e^2\] \(\large e^{ln} = 1\) so we have \[\large x - 2 = e^2\] now you would just solve for 'x'

OpenStudy (anonymous):

so I would basically just solve x-2=2 fir x

OpenStudy (anonymous):

for

OpenStudy (johnweldon1993):

Well \[\large x - 2 = e^2\] for 'x' yes (you forgot the 'e' above\]

OpenStudy (anonymous):

so add 2 right

OpenStudy (johnweldon1993):

Correct

OpenStudy (anonymous):

but then how would I solve x=2e^2

OpenStudy (anonymous):

put it in my calculator

OpenStudy (johnweldon1993):

you mean x = e^2 + 2 right?

OpenStudy (anonymous):

ohhhh ok sry my bad

OpenStudy (anonymous):

so 2.6931 = x

OpenStudy (johnweldon1993):

Not quite what I get....was that out of your calculator? Well put it like this... e is approximately equal to 2.718 (If I remember correctly so we have \[\large x = 2.718^{2} + 2\] what does that come out to?

OpenStudy (anonymous):

9.3875 and I would plug 2.718 to any problem similar to this one

OpenStudy (johnweldon1993):

Well that is just approximate...if you want the exact answer..use the 'e' key in your calculator... But yes you can do that

OpenStudy (anonymous):

ok and umm do you mind helping me with one more problem.